OpenStudy (gabylovesyou):

Which mathematical sentence represents the solution for d in the equation 6e = ef + 3d? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

what do you get when you solve for d

jimthompson5910 (jim_thompson5910):

If you're stuck, then I'll start you off \[\large 6e = ef + 3d\] \[\large 6e - ef = 3d\] What comes next?

OpenStudy (gabylovesyou):

um add ef ?

jimthompson5910 (jim_thompson5910):

that will just get you back to \[\large 6e = ef + 3d\]

jimthompson5910 (jim_thompson5910):

you want to isolate d

jimthompson5910 (jim_thompson5910):

3d means 3 times d how do you undo the multiplication of 3?

OpenStudy (gabylovesyou):

dividing? ;/

jimthompson5910 (jim_thompson5910):

yep, divide both sides by what

OpenStudy (gabylovesyou):

3 ?

jimthompson5910 (jim_thompson5910):

so in full, your steps will look like this \[\large 6e = ef + 3d\] \[\large 6e - ef = 3d\] \[\large 3d = 6e - ef\] \[\large d = \frac{6e - ef}{3}\]

jimthompson5910 (jim_thompson5910):

first you subtract ef from both sides then you divide both sides by 3 to fully isolate d.

OpenStudy (gabylovesyou):

ohh ok :)